OpenStudy (anonymous):

Simplify x^6 times x^-4

OpenStudy (anonymous):

i got 100

OpenStudy (sidsiddhartha):

use\[(x ^{a}*x ^{b})=x ^{a+b}\]

OpenStudy (anonymous):

x^2 1/x^2 x^10 1/x^10

OpenStudy (solomonzelman):

\[x^{6} \times x^{-4}=x^{6+(-4)}=x^{?}\]

OpenStudy (anonymous):

would it be x^2

OpenStudy (solomonzelman):

yes

OpenStudy (sidsiddhartha):

yeah

OpenStudy (anonymous):

thanks. Can you help me with another one?

OpenStudy (sidsiddhartha):

yeah

OpenStudy (anonymous):

Thanks. simplify. (x^4)^2 times x^3

OpenStudy (anonymous):

woud it be x^11? @sidsiddhartha

OpenStudy (solomonzelman):

yes

OpenStudy (sidsiddhartha):

yeah correct:)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

simplify 6 - 5(3 - 4x) (this one is a little confusing)

OpenStudy (sidsiddhartha):

\[6-5(3-4x)=6-(5*3)-(5*-4x)=6-15+20x\]

OpenStudy (anonymous):

9-4x 6-5x 3-4x 20x-9 would it be this one?